OpenStudy (anonymous):

Does 5xy+y/2x Divided By y/6x equal 3(5x+1)?

OpenStudy (anonymous):

\[\frac{\frac{5xy+y}{2x}}{\frac{y}{6}}\] like this ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[\frac{\frac{5xy+y}{2x}}{\frac{y}{6x}}\] like this ?

OpenStudy (anonymous):

except it is horizontal, but that' like the same thing

OpenStudy (anonymous):

that's*

OpenStudy (anonymous):

yeaah your answer is correct ! let me check again !

OpenStudy (anonymous):

okay :) thanks

OpenStudy (anonymous):

yeaaah It's 3(5x+1)

OpenStudy (anonymous):

:) you're welcome anytime !